OpenStudy (anonymous):

A line has a slope of 1 and passes through the point (6, 3). What is its equation in slope-intercept form?

terenzreignz (terenzreignz):

Well... if you have a slope m and the line passes through the point (a,b) then the line has the equation... \[\huge y = \color{blue}m(x-a) + b\]

OpenStudy (anonymous):

Then you would have y=1x-6+3

terenzreignz (terenzreignz):

Actually, yes :) You can do away with the 1... And you'll have...?

OpenStudy (anonymous):

y=x-6+3

terenzreignz (terenzreignz):

Come on, you can still do stuff with that... \[\large y = x-6+3\\ \large y = x \boxed{\color{red}{-6+3}}\]

OpenStudy (anonymous):

y=x-3

terenzreignz (terenzreignz):

Yes, perfect :)

OpenStudy (anonymous):

Thank you so much:)

terenzreignz (terenzreignz):

No problem :) I'm TJ, welcome to OpenStudy... Stick around, see if you can help with some other guys with questions :D

OpenStudy (anonymous):

okay one last question The points (0, 8) and (-20, 12) fall on a particular line. What is its equation in slope-intercept form?

terenzreignz (terenzreignz):

Okay, you need a slope here... If you have two points, (a,b) and (c,d) the slope is given by this formula... \[\Large m = \frac{d-b}{c-a}\] So, can you get the slope?

OpenStudy (anonymous):

1/-5

terenzreignz (terenzreignz):

okay... \[\Large m = -\frac15\] That's very good... Now you use the formula again... we have a slope now... \[\LARGE y = m(x-a) + b\] As for your point, you can use any one of the two points you have been given, it doesn't matter...

OpenStudy (anonymous):

okay so \[y=1/-5(x-0)+8\]

terenzreignz (terenzreignz):

Yeah, careful...I know you have the correct idea... let me just rewrite it for you... \[\Large y = -\frac15(x-0) + 8\]

OpenStudy (anonymous):

So the final product would be \[y=-1/5x-8\]

OpenStudy (anonymous):

I mean +8

terenzreignz (terenzreignz):

Okay, I just hope you meant this... \[\LARGE y = -\frac15x + 8 \] It seems you have it down, good job :)